# Are We There Yet? A Brief Survey of Music Emotion Prediction Datasets, Models and Outstanding Challenges

### v2

@@ Abstract
Deep learning models for music have advanced drastically in recent years, but how good are machine learning models at capturing emotion, and what challenges are researchers facing? In this paper, we provide a comprehensive overview of the available music-emotion datasets and discuss evaluation standards as well as competitions in the field. We also offer a brief overview of various types of music emotion prediction models that have been built over the years, providing insights into the diverse approaches within the field. Through this examination, we highlight the challenges that persist in accurately capturing emotion in music, including issues related to dataset quality, annotation consistency, and model generalization. Additionally, we explore the impact of different modalities, such as audio, MIDI, and physiological signals, on the effectiveness of emotion prediction models. Recognizing the dynamic nature of this field, we have complemented our findings with an accompanying GitHub repository111 https://github.com/AMAAI-Lab/awesome-MER/. This repository contains a comprehensive list of music emotion datasets and recent predictive models.
@@ I Introduction
Despite these challenges, the potential applications of Music Emotion Recognition (MER) systems are vast and varied. In the healthcare domain, we may find personalized music recommendation systems that guide a listener to different emotional states [5], or we may even see MER systems used to build large-scale datasets on which we can train generative music AI systems that can be controlled to generate music with specific emotions [6, 7, dashherremans2017morpheus]. In the same line of thought, [8] developed a Brain-Computer Interface system that can provide musical feedback about the listener’s current emotional state and subsequently influence this state. Additionally, MER systems may help facilitate emotional analysis during music composition, interactive experiences in media and entertainment, as well as inform market research [9]. The implications of understanding music and emotions extend across diverse domains. In general, following the idea of positive psychology [10], understanding music emotions can be used to enhance the user experience when designing various systems.
@@
YM2413-MDB [47] 2022 699 full WAV, MIDI 35 labels - S P
@@ II Datasets
Table I presents an extensive overview of emotion-annotated music datasets. These datasets vary in size, annotation granularity, and focus on either perceived or induced emotions. We have attempted to provide an exhaustive overview of emotion datasets. This was achieved by using Google Scholar with search terms such as ‘music emotion dataset’, ‘affective music dataset’, as well as following references within these articles. Before delving deeper into the datasets, it is essential to understand the emotion representations commonly referenced in the field.
@@
Emotion representations One of the earliest emotion representation models in music is Hevner’s affective ring [58], developed in 1936. Based on extensive experimental studies, Hevner’s model categorizes music emotions into eight fundamental categories: dignified, sad, dreamy, serene, graceful, happy, exciting, and vigorous. In current-day MER research, we see that Russell’s Circumplex Model of Affect [59] is widely used. This model characterizes emotions along two dimensions: valence and arousal. Valence represents the degree of positive or negative emotion, while arousal reflects the intensity of emotion, ranging from passive to activated states. Russell’s original model contained a third dimension: dominance [59]. This dimension is typically omitted as it can be hard to annotate, although some researchers have argued to re-include it [60]. Russell’s model is a dimensional model, as it consists of continuous values along multiple dimensions (valence/arousal). Hevner’s model, on the other hand, is categorical as it consists of discrete emotion labels.
Thayer’s two-dimensional model [61] focuses on energetic arousal and tense arousal as the primary dimensions of emotion. Thayer suggests that valence can be inferred from the combination of energetic and tense arousal levels.
@@
Categorical models include the Geneva Emotional Music Scales (GEMS) [27]. This model was specifically designed for music-induced emotions and consists of 45 emotion tags grouped into nine categories, including amazement, solemnity, tenderness, nostalgia, calmness, power, joyful activation, tension, and sadness. The datasets listed in Table I use a variety of emotion representation models, with Russell’s model being the most popular dimensional representation. Some datasets do not use a specific emotion model, for instance, the MTG-Jamendo dataset [36] consists of freely assigned tags by the listeners, resulting in a diverse and comprehensive set of 56 tags, spanning from ‘melancholic’ to ‘upbeat’. A subset of this dataset is used for the ‘Emotion and Theme Recognition in Music’ Task of MediaEval. In Table I, the number of tags used to represent the emotions are listed in the column ‘Categorical’.
@@
Annotation Process Annotation processes and label distribution vary significantly across datasets, impacting the effectiveness and reliability of machine learning models in MER. Some datasets, like MTG-Jamendo [36], rely on freely assigned tags by listeners, allowing for a broad range of emotional descriptors. In contrast, DEAP [22] and HKU956 [44] use a predetermined list of emotions to ensure consistency in emotion annotations. The number of annotators also differs, with some datasets involving large crowdsourced annotations, providing diverse perspectives, while others use smaller, more controlled groups to maintain uniformity. Additionally, the logic behind selecting music segments varies, with some datasets focusing on full-length songs and others on shorter clips to capture transient emotional responses.
@@
Label distribution and splits The distribution of labels across datasets can be imbalanced, affecting the performance of machine learning models. For instance, the MTG-Jamendo dataset [36] contains a diverse set of 56 tags, but the frequency of these tags can vary significantly. In contrast, the DEAP dataset [22] provides a more balanced set of labels across arousal and valence dimensions. Finally, some datasets, such as DEAM [32], propose fixed splits for training and testing. This facilitates the replication of experiments and ensures consistent evaluation metrics which are easily used as a benchmark.
@@ III Evaluation Protocols
Given the noisiness of labels in music-emotion datasets [45], one could also argue to include listening tests to verify the accuracy of the predicted labels, or even to perform extrinsic evaluation through a downstream task. The latter could be implementing MER in an emotion-inducing system, e.g. by performing playlist recommendations based on predicted emotions from music in the dataset, and subsequently evaluating if the listeners’ emotion indeed reaches the target emotion.
@@
TABLE II: Overview of Selected Music Emotion Recognition Models since 2020.
Ref. Year Modalities Approach Emotion Model Dataset
[76] 2020 Audio, Lyrics Machine learning (e.g., SVM, NB) Russell PMEmo
[77] 2020 Audio CNN, LSTM Russell Solymani et al.’s dataset
@@
[78] 2020 Audio LSTM 4 classes Self-built
[79] 2020 Audio CNN-LSTM 4 classes Last.fm
[80] 2020 Audio Attentive LSTM Russell Solymani et al.’s dataset
[81] 2020 Audio Source separation, CNN Russell PMEmo
[82] 2020 Audio Cochleogram, CNN Russell Solymani et al.’s dataset
[83] 2020 Audio CNN, Local Attention Quadrants Soundtrack, Bi-Modal
@@
[72] 2020 Audio Attention Based Neural Networks 56 classes MTG-Jamendo
@@
[71] 2020 Audio Triplet Neural Networks Russell DEAM
[84] 2021 Audio CNN-LSTM Russell DEAM
@@
[85] 2021 Audio CNN, LSTM+DNN 3 classes Self-built
@@
[73] 2021 Audio Clustering-based Ensembles 56 classes MTG-Jamendo
@@
[74] 2021 Audio Semi-Supervised, Noisy Student Training 56 classes MTG-Jamendo
@@
[75] 2021 Audio Frequency Dependent Convolutions 56 classes MTG-Jamendo
@@
[86] 2021 Audio CNN, Co-teaching Training Strategy 56 classes MTG-Jamendo
[87] 2021 Audio LSTM, Pretrained Models Russell Self-built
[88] 2021 Lyrics Transformers Russell MoodyLyrics, MER
@@
[89] 2021 Audio Source-separation based explainer Russell DEAM, Midlevel, PMEmo
@@
[90] 2021 Audio Generative Adversarial Network 2 classes Self-built
[91] 2021 Audio, Video CNN 6 classes Self-built
[92] 2021 Audio Linear Regressors Russell Self-built
@@
[93] 2022 Audio Clustering, Machine Learning (e.g., SVM) Russell Solymani et al.’s dataset
@@
[94] 2022 MIDI Multi-Task Learning Quadrants EMOPIA, VGMIDI
[95] 2022 Audio Feature Selection, SVR, RF Russell DEAM
[46] 2022 Audio, Video LSTM Russell Self-built
[96] 2022 Lyrics Deep Learning, BERT Quadrants MIR Lyrics Emotion
[97] 2022 Audio CNN, LSTM Russell PMEmo
[98] 2022 Audio SVM, Random Forest, MLP Russell DEAM
[99] 2023 Audio, Lyrics Multi-Modality Russell DMDD
[100] 2023 Audio Inception-GRU Residual Quadrants Soundtrack
[101] 2023 Lyrics State Space Models 6 classes LyricsEmotions
@@
[102] 2023 Lyrics Fine-tuned XLMRoBERTa Russell Self-built
@@
[50] 2023 Audio, Lyrics CNN-LSTM, XLNet Transformers 3 classes Self-built
[103] 2023 Audio Attention Mechanism Russell DEAM, PMEmo
[104] 2023 Audio CNN 4 classes Musical Excerpts
[105] 2023 Audio, Lyrics Transformers Quadrants MoodyLyrics
[106] 2024 Audio RNN, BRNN, LSTM Quadrants 4Q audio, MTG-Jamendo
@@
[107] 2024 Audio, Lyrics VGGish, ALBERT 4 classes DEAM, FMA
@@
[108] 2024 Audio, Lyrics Stacked Ensemble Models 4 classes Self-built
@@
[109] 2024 Audio Multi-scale Parallel Convolution 8 classes PMEmo, Soundtrack, RAVDESS
@@
[110] 2024 Audio CNN with Differential Evolution 4 classes self-built, DEAM

### v3

@@ Abstract
Deep learning models for music have advanced drastically in recent years, but how good are machine learning models at capturing emotion, and what challenges are researchers facing? In this paper, we provide a comprehensive overview of the available music-emotion datasets and discuss evaluation standards as well as competitions in the field. We also offer a brief overview of various types of music emotion prediction models that have been built over the years, providing insights into the diverse approaches within the field. Through this examination, we highlight the challenges that persist in accurately capturing emotion in music, including issues related to dataset quality, annotation consistency, and model generalization. Additionally, we explore the impact of different modalities, such as audio, MIDI, and physiological signals, on the effectiveness of emotion prediction models. Through this examination, we identify persistent challenges in music emotion recognition (MER), including issues related to dataset quality, the ambiguity in emotion labels, and the difficulties of cross-dataset generalization. We argue that future advancements in MER require standardized benchmarks, larger and more diverse datasets, and improved model interpretability. Recognizing the dynamic nature of this field, we have complemented our findings with an accompanying GitHub repository111 https://github.com/AMAAI-Lab/awesome-MER/. This repository contains a comprehensive list of music emotion datasets and recent predictive models.
@@ I Introduction
Despite these challenges, the potential applications of Music Emotion Recognition (MER) systems are vast and varied. In the healthcare domain, we may find personalized music recommendation systems that guide a listener to different emotional states [5], or we may even see MER systems used to build large-scale datasets on which we can train generative music AI systems that can be controlled to generate music with specific emotions [6, 7, 8]. In the same line of thought, [9] developed a Brain-Computer Interface system that can provide musical feedback about the listener’s current emotional state and subsequently influence this state. Additionally, MER systems may help facilitate emotional analysis during music composition, interactive experiences in media and entertainment, as well as inform market research [10]. The implications of understanding music and emotions extend across diverse domains. In general, following the idea of positive psychology [11], understanding music emotions can be used to enhance the user experience when designing various systems.
@@
Earlier surveys on Music Emotion Recognition (MER) include the review by Panda et al. (2020) [16], which focuses on the taxonomy of audio features and their relationship to emotion perception. More recently, a broad overview of deep learning techniques for MER was provided in [17], with extended discussion on emotion representations. However, these surveys primarily focus on either specific feature engineering strategies or model architectures, and they often lack comprehensive coverage of datasets, evaluation protocols, and the challenges of integrating multimodal sources or predicting induced emotion. In contrast, our goal is to provide a broader and more integrative perspective on the MER landscape. We examine not only the evolution of model architectures, but also the diversity of datasets, annotation strategies, emotion models, and evaluation practices. This comprehensive perspective allows us to highlight emerging trends, identify gaps in current methodologies, and suggest promising directions for future research.
@@
Last.fm subset of MSD [25] 2011 505,216 full Metadata only listener tags - S P
@@
YM2413-MDB [52] 2022 699 full WAV, MIDI 19 labels - S P
@@
Affolter and Rohrmeier’s dataset [63] 2024 5,892 full Spotify id 8 labels - S P
XMIDI [64] 2025 108,023 full MIDI 11 labels - S P
@@ II Datasets
Table I presents an extensive overview of emotion-annotated music datasets. These datasets vary in size, annotation granularity, and focus on either perceived or induced emotions. We have attempted to provide an exhaustive overview of emotion datasets. This was achieved by using Google Scholar with search terms such as ‘music emotion dataset’, ‘affective music dataset’, as well as following references within these articles. Before delving deeper into the datasets, we first describe key dimensions that help structure our comparison. These include: (1) the emotion representation model (e.g., categorical vs. dimensional), (2) the annotation method (static or dynamic), (3) the type of emotion captured (perceived vs. induced), (4) the data modalities involved (e.g., audio, video, lyrics), and (5) the dataset’s size and diversity in genre or listener demographics. Below, we elaborate on each of these aspects and illustrate them with representative examples.
@@
Emotion representations One of the earliest emotion representation models in music is Hevner’s affective ring [65], developed in 1936. Based on extensive experimental studies, Hevner’s model categorizes music emotions into eight fundamental categories: dignified, sad, dreamy, serene, graceful, happy, exciting, and vigorous. In current-day MER research, we see that Russell’s Circumplex Model of Affect [66] is widely used. This model characterizes emotions along two dimensions: valence and arousal. Valence represents the degree of positive or negative emotion, while arousal reflects the intensity of emotion, ranging from passive to activated states. Russell’s original model contained a third dimension: dominance [66]. This dimension is typically omitted as it can be hard to annotate, although some researchers have argued to re-include it [67]. It is worth noting, however, that the dominance dimension was more central to Russell’s earlier PAD (Pleasure-Arousal-Dominance) framework, designed for environmental and contextual emotion analysis. Its omission from the Circumplex Model was a simplification intended to facilitate self-reported emotion studies.
Russell’s model is a dimensional model, as it consists of continuous values along multiple dimensions (valence/arousal). Hevner’s model, on the other hand, is categorical as it consists of discrete emotion labels. Thayer’s two-dimensional model [68] focuses on energetic arousal and tense arousal as the primary dimensions of emotion. Thayer suggests that valence can be inferred from the combination of energetic and tense arousal levels. Other categorical models include the Geneva Emotional Music Scales (GEMS) [31]. This model was specifically designed for music-induced emotions and consists of 45 emotion tags grouped into nine categories, including amazement, solemnity, tenderness, nostalgia, calmness, power, joyful activation, tension, and sadness.
Recently, alternative emotion frameworks grounded in psychological theory have also been explored. For instance, Affolter et al. [63] introduced a dataset that maps listener-generated tags and playlist names to emotion labels derived from Plutchik’s psychoevolutionary model of emotions. Their method employs natural language processing to associate each track with an 8-dimensional vector corresponding to Plutchik’s basic emotions, offering a novel bridge between tag-based annotations and theoretical models of affect.
@@
The datasets listed in Table I use a variety of emotion representation models, with Russell’s model being the most popular dimensional representation. Some datasets do not use a specific emotion model, for instance, the MTG-Jamendo dataset [41] consists of freely assigned tags by the listeners, resulting in a diverse and comprehensive set of 56 tags, spanning from melancholic’ to upbeat’. A subset of this dataset is used for the Emotion and Theme Recognition in Music’ Task of MediaEval [69], which serves as a benchmark for evaluating MER systems. In Table I, the number of tags used to represent the emotions are listed in the column “Categorical”.
@@
Annotation Process Annotation processes and label distributions vary significantly across datasets, impacting the effectiveness and reliability of machine learning models in MER. Some datasets, such as MTG-Jamendo [41], rely on freely assigned tags by multiple annotators, allowing a broad range of emotional descriptors. In contrast, DEAP [26] and HKU956 [49] use a predetermined list of emotions to ensure consistency across annotations.
A major distinction between annotation strategies relates to if they use absolute or relative annotation methods, this is particularly important to note for dimensional models of emotion. In absolute annotation, annotators assign a direct score or category to each sample independently. This approach is common in datasets such as DEAP and HKU956, where participants provide valence and arousal ratings individually for each stimulus. Recenlty, several datasets propose relative annotation strategies to simplify the annotation process and improve consistency across annotators. Instead of rating individual tracks absolutely, annotators are asked to compare pairs of samples and judge their relative emotional positioning. Examples include the MusAV dataset [53], which gathers comparative annotations of arousal and valence for pairs of music tracks, and Emo-Soundscapes [78], where participants rank soundscape recordings based on perceived emotion. Earlier work by Yang et al. [79] introduced ranking-based emotion recognition methods for music organization and retrieval, showing that relative judgments can lower cognitive load and lead to more reliable ground truth data. Similarly, in the EMusic dataset [80], a ranking-based annotation approach was used for experimental music, demonstrating improvements over traditional absolute labeling methods.
In addition, the CCMED-WCMED dataset [47] compared Western and Chinese classical music based on emotion dimensions collected through relative annotations, highlighting the importance of cultural factors in emotion perception. These relative approaches can mitigate some of the variability and subjectivity inherent in absolute annotations, particularly in continuous emotion models like valence-arousal space.
The number of annotators also varies: some datasets involve large-scale crowdsourcing efforts, providing a wide range of perspectives (e.g., MusAV, Emo-Soundscapes), while others use smaller, more controlled groups to maintain annotation consistency (e.g., DEAP, HKU956). Additionally, the logic behind selecting music segments differs, with some datasets focusing on full-length songs and others on shorter clips to better capture transient emotional responses.
Overall, the choice between absolute and relative annotation methods, the number of annotators, and the segment selection strategies are all critical factors influencing dataset reliability and the resulting model performance in MER research.
Figure 1: Top 10 most and least frequent tags in the MTG-Jamendo dataset.
@@
Label distribution and splits The distribution of labels across datasets can be imbalanced, affecting the performance of machine learning models. For instance, the MTG-Jamendo dataset [41] contains a diverse set of 56 tags, freely assigned by users, but their frequency is highly imbalanced. This skewed distribution, with some tags like “happy” and “film” appearing far more often than others like “sexy” or “fast”, is visualized in Figure 2, which shows the top 10 most and least frequent tags. In contrast, the DEAP dataset [26] provides a more balanced set of labels across arousal and valence dimensions. Finally, some datasets, such as DEAM [37] and MTG-Jamendo [41], propose fixed splits for training and testing. This facilitates the replication of experiments and ensures consistent evaluation metrics which are easily used as a benchmark.
Recommendations and Future Directions While the landscape of emotion-annotated music datasets is expanding, several challenges and opportunities remain. For general-purpose MER models, datasets like MTG-Jamendo [41] and MERP [50] offer breadth in genre and label diversity, while EMOPIA [44] and VGMIDI [42] are more suitable for symbolic-domain or piano-focused emotion modeling. For multimodal applications, Popular Hooks [62] and EmoMV [54] are relevant. However, there is a notable lack of large-scale, publicly available datasets focusing on induced emotions with multimodal or biosensor data, limiting medical and affective computing applications. The community would benefit from more datasets capturing diverse demographics, cross-cultural emotional interpretations, and consistent emotion annotation schemas across modalities. We encourage future dataset creators to include dynamic annotations, listener metadata, and align modalities when possible.
@@ III Evaluation Protocols
The evaluation approach also depends on whether a dataset provides static or dynamic annotations. Static annotations assign a single label (or vector) to an entire track or segment, making evaluation straightforward with standard classification or regression metrics (e.g., accuracy, F1-score, ROC-AUC, or $R^{2}$) computed at the song level [41, 36, 20]. In contrast, dynamic annotations provide time-continuous emotion ratings—often sampled at a rate of 1 Hz—reflecting how emotions evolve throughout a track [37, 26, 38]. In such cases, evaluation typically involves comparing predicted and ground-truth time series using frame-level metrics such as Pearson Correlation Coefficient (PCC), Mean Squared Error (MSE), or Concordance Correlation Coefficient (CCC) [37, 82, 38]. Some studies apply smoothing or window-based aggregation (e.g., moving average or median filters) to reduce annotation noise and capture broader emotional trends [82, 37]. For example, the DEAM dataset [37] provides both static and dynamic valence-arousal annotations, with dynamic evaluation conducted on a per-frame basis across time.
When evaluating MER models, it is also important to distinguish between perceived and induced (invoked) emotion. Evaluating perceived emotion predictions typically involves comparing predicted labels with human-annotated ground truth, which reflects how listeners interpret the emotional content of the music. In contrast, evaluating induced emotion is more complex, as it involves the actual emotional response elicited in the listener. This often requires physiological measurements (e.g., heart rate, EEG) or self-reports in a controlled experimental setup. For this reason, evaluation of induced emotion prediction models may rely on additional modalities (e.g., biosensors), and often focuses on subjective measures such as emotion regulation effectiveness or user satisfaction. These differences highlight the need for tailored evaluation protocols depending on the target emotion type.
@@
Given the variability and potential noisiness of emotion annotations in music datasets [50], evaluation protocols should be designed to assess not only performance against possibly imperfect ground truth but also the perceptual validity of model outputs. One way to address this is by incorporating listening tests, where human subjects rate the emotions conveyed by model-predicted tracks. This does not directly mitigate the noise in the training data, but rather provides an additional, human-centered validation of the model’s effectiveness. Another complementary evaluation strategy is extrinsic evaluation, where emotion recognition is applied in downstream tasks. For instance, predicted emotions could be used to generate emotionally coherent playlists, and users could be asked whether the playlists evoke the intended emotional responses. Such evaluations help assess whether MER systems produce outputs that are meaningful and useful in real-world applications, beyond what can be inferred from standard metrics alone.
@@
TABLE II: Overview of Selected Music Emotion Recognition Models since 2020. Note: V = Valence, A = Arousal, Q = Quadrant.
Ref. Year Modalities Approach Emotion Model Dataset Performance
[87] 2020 Audio, Lyrics Machine Learning (e.g., SVM, NB) Russell PMEmo Accuracy: 63%
[88] 2020 Audio CNN, LSTM Russell Solymani et al.’s dataset RMSE: 0.219 V, 0.212 A
@@
[89] 2020 Audio LSTM happy, sad, neutral, fear Self-built Accuracy: 89.3%
[90] 2020 Audio, Lyrics CNN-LSTM angry, happy, relaxed, sad Last.fm Accuracy: 78%
[91] 2020 Audio Attentive LSTM Russell Solymani et al.’s dataset R2: 0.53 V, 0.75 A
[92] 2020 Audio Source Separation, CNN Russell PMEmo R2: 0.4814 V, 0.6004 A
[93] 2020 Audio Cochleogram, CNN Russell Solymani et al.’s dataset R2: 0.41 V, 0.63 A
[94] 2020 Audio CNN, Local Attention Quadrants Soundtrack, Bi-Modal Accuracy (Soundtrack): 67.71%
F1 (Bi-Modal): 77.82%
@@
[83] 2020 Audio Attention-based Neural Networks 56 mood/theme tags MTG-Jamendo PR-AUC: 0.118, ROC-AUC: 0.735
@@
[82] 2020 Audio Triplet Neural Networks Russell Solymani et al.’s dataset, DEAM R2 (Solymani.): 0.378 V, 0.638 A
R2 (DEAM): 0.361 V, 0.672 A
[95] 2020 Audio CNN Ensemble, Focal Loss, Receptive Field Tuning 56 mood/theme tags MTG-Jamendo PR-AUC: 0.161, ROC-AUC: 0.781
[96] 2021 Audio CNN-LSTM Russell DEAM Accuracy: 64.9%
@@
[97] 2021 Audio CNN, LSTM+DNN 3 classes (high arousal + pos./neg. valence) Self-built Accuracy: 99.19%
@@
[84] 2021 Audio Clustering-based Ensembles 56 mood/theme tags MTG-Jamendo PR-AUC: 0.109, ROC-AUC: 0.705
@@
[85] 2021 Audio Semi-Supervised, Noisy Student Training 56 mood/theme tags MTG-Jamendo PR-AUC: 0.136, ROC-AUC: 0.769
@@
[86] 2021 Audio Frequency Dependent Convolutions 56 mood/theme tags MTG-Jamendo PR-AUC: 0.151, ROC-AUC: 0.775
@@
[98] 2021 Audio CNN, Co-teaching Training Strategy 56 mood/theme tags MTG-Jamendo PR-AUC: 0.144, ROC-AUC: 0.760
[99] 2021 Audio LSTM, Pretrained Models Russell Self-built R2: 0.46 V, 0.73 A
[100] 2021 Lyrics Transformers Russell MoodyLyrics, MER Accuracy (MoodyLyrics): 94.78%
Accuracy (MER): 94.44% V, 88.89% A,
88.89% Q
@@
[101] 2021 Audio Source-separation-based Explainer Russell DEAM, Midlevel, PMEmo R2 (DEAM): 0.48 V, 0.50 A
R2 (PMEmo): 0.50 V, 0.65 A
@@
[102] 2021 Audio Generative Adversarial Network happy, sad Self-built Accuracy: 87%
[103] 2021 Audio, Video CNN exciting, fear, neutral, relaxation, sad, tension Self-built Accuracy: 88.56%
[104] 2021 Audio Linear Regressors Russell Self-built R2: 0.78 V, 0.85 A
@@
[105] 2022 Audio Clustering, Machine Learning (e.g., SVM) Russell Solymani et al.’s dataset Accuracy: 84.9%
@@
[106] 2022 MIDI Multi-Task Learning Quadrants EMOPIA, VGMIDI Accuracy (EMOPIA): 67.58%
Accuracy (VGMIDI): 55.85%
[107] 2022 Audio Feature Selection, SVR, RF Russell DEAM R2: 0.587 V, 0.645 A
[51] 2022 Audio, Video LSTM Russell Self-built RMSE: 0.1331 V, 0.0973 A
[108] 2022 Lyrics Deep Learning, BERT Quadrants MIR Lyrics Emotion F1: 88.9 %
[109] 2022 Audio CNN, LSTM Russell PMEmo, AllMusic Accuracy (PMEmo):
79.01% V, 83.62% A
Accuracy (AllMusic):
67.11% V, 86.56% A
[110] 2022 Audio SVM, Random Forest, MLP Russell DEAM RMSE: 0.23 V, 0.24 A
[111] 2023 Audio, Lyrics Multi-Modality Russell DMDD R2: 0.235 V, 0.196 A
[112] 2023 Audio Inception-GRU Residual Quadrants Soundtrack Accuracy: 84%
[113] 2023 Lyrics State Space Models anger, disgust, fear, joy, sadness, surprise LyricsEmotions Pearson correlation: 0.345 (angry),
0.268 (disgust), 0.350 (fear),
0.503 (joy), 0.350 (sad),
0.089 (surprise)
@@
[114] 2023 Lyrics Fine-tuned XLMRoBERTa Russell Self-built Accuracy: 80.88% V, 81.51% A,
62.58% Q
@@
[55] 2023 Audio, Lyrics CNN-LSTM, XLNet Transformers positive, neutral, negative Self-built Accuracy: 80.56%
[115] 2023 Audio Attention Mechanism Russell DEAM, PMEmo RMSE (DEAM): 0.112 V, 0.109 A
RMSE (PMEmo): 0.144 V, 0.135 A
[116] 2023 Audio CNN happiness, fear, sadness, peacefulness Musical Excerpts Accuracy: 92%
[117] 2023 Audio, Lyrics Transformers Quadrants MoodyLyrics Accuracy: 77.94%
[118] 2024 Audio RNN, BRNN, LSTM Quadrants 4Q audio, MTG-Jamendo Accuracy (4Q.): 65.97%
Accuracy (MTG.): 53.97%
@@
[119] 2024 Audio, Lyrics VGGish, ALBERT happy, sad, calm, healing DEAM, FMA Accuracy (DEAM): 49.68%
Accuracy (FMA): 49.54%
TABLE II: (continued)
Ref. Year Modalities Approach Emotion Model Dataset Performance
@@
[120] 2024 Audio, Lyrics Stacked Ensemble Models angry, happy, relax, sad Self-built Accuracy: 96.25%
@@
[121] 2024 Audio Multi-scale Parallel Convolution sleepy, calm, sad, pleased, relaxed, nervous, annoying, excited PMEmo, Soundtrack, RAVDESS Accuracy: 98.58%
@@
[122] 2024 Audio CNN with Differential Evolution happy, sad, angry, calm self-built, DEAM Accuracy: 85.29%
[123] 2024 Audio LLM Embeddings, Non-parametric Clustering Categorial MTG-Jamendo, CAL500, Emotify F1 (MTG-Jamendo): 2.02%
F1 (CAL500): 22.9%
F1 (Emotify): 40.0%
[124] 2025 Audio MERT Embeddings, Multitask Learning, Transformers, Knowledge Distillation Categorial, Russell MTG-Jamendo, PMEmo, DEAM, Solymani et al.’s dataset (MTG-Jamendo) PR-AUC: 0.1543
ROC-AUC: 0.7810
(Solymani.) R2: 0.6512 V, 0.7616 A
(PMEmo) R2: 0.5473 V, 0.7940 A
(DEAM) R2: 0.5184 V, 0.6228 A